What do you understand by half-wave rectification of half-wave rectification of an alternating current? Obtain the following values of half-wave rectified A.C.

a)     RMS value

b)     Average value 


1
Expert's answer
2021-08-31T14:22:39-0400

The half-wave rectification of half-wave rectification of an alternating current is a type of rectification method that, receiving a sine or cosine wave on the input, outputs only upper or lower part of the wave, i.e. above or below the x-axis.

The RMS value:


"V_\text{rms}=\frac{V_\text{max}}{2}."

The average value:


"V_\text{avg}=\frac{V_\text{max}}{\pi}."